# MSTAR: Box-free Multi-query Scene Text Retrieval with Attention Recycling

### v1

@@ B.2 Visualization of the MQTR dataset
To clarify the characteristics of the MQTR dataset, we show some examples collected from the google image search engine, as presented in Tab.  10. Since scene text typically appears as finegrained information but the search engine often recommends the most salient images, we collect more images containing queries in fine-grained concepts. For example, “global weekly” is shown as a section name of “china daily” newspaper, adding more challenges for models. For example, for the query “dream big”, the model may have difficulty distinguishing between “bream big” written on the t-shirt and “dream big” written on the shorts.
@@ Appendix C Experiment Details and Visualization Analysis
C.1 Training Details
To ensure the reproducibility of our model, we provide the training process and the training parameters, which are reported in Tab. 11. We adopt a progressive training recipe. In the first state, the model is trained on both synthetic data $D_{\text{syn}}$ and real data $D_{\text{real}}$ at an image resolution of 512. We use a learning rate of 1e-5 and linear cosine schedule with 100 warmming-up steps. In this stage, we use a two-layer MLP to align the visual encoder and multi-modal encoder. In the second stage, our model is finetuned at the resolution of 640 only on the real data $D_{\text{real}}$. For simplicity, the hyperparameters keep the same with the first stage. In the third stage, we finetune the model at a higher resolution of 800 on the real data $D_{\text{real}}$. In the forth stage, the visual encoder is freezed and only the MLP, multi-modal encoder, MIM, and SAS module are optimized at a resolution of 800. For the word retrieval experiment, the synthetic data $D_{\text{syn}}$ refers to 100K images randomly sampled from SynthText-900k, and the real data $D_{\text{real}}$ is MLT-5K. For the multi-query experiment, $D_{\text{syn}}$ includes 50K images randomly sampled from SynthText-900k and 25k images with phrase transcriptions with the synthesis engine. $D_{\text{real}}$ is the training set from the TextCap dataset. The labels are the image captions and text transcriptions acquired with Rosetta.
C.2 Experimental Visualization Analysis
We present a qualitative analysis of the retrieval results. As illustrated in Tab.  12, our method can not only effectively leverage linguistic priors for phrase and semantic queries, but also perceive fine-grained scene text instances to achieve word and combined retrieval. For example, MSTAR successfully retrieves the combined query like “reserved” and “some” even when they appear as subtle watermarks in pictures, as depicted in the last row.
@@ C.2 Experimental Visualization Analysis
“speed limit 25” Bridge  [11] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
“hydrate or diedrate” is written on a water bottle Bridge  [11] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
“may” BLIP2 (FT)  [19] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
“reserved” , “some” BLIP2 (FT) [19] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 12: Qualitive analysis of the retrieval results. The green boxes mean correct samples and the red boxes denote false images. The queries and images are sampled from the MQTR dataset and CTW [21] dataset.

### v2

@@ B.2 Visualization of the MQTR dataset
To clarify the characteristics of the MQTR dataset, we show some examples collected from the google image search engine, as presented in Tab.  10. Since scene text typically appears as fine-grained information but the search engine often recommends the most salient images, we collect more images containing queries in fine-grained concepts. For example, “global weekly” is shown as a section name of “china daily” newspaper, adding more challenges for models. For example, for the query “dream big”, the model may have difficulty distinguishing between “bream big” written on the t-shirt and “dream big” written on the shorts.
B.3 Statistics of the MQTR dataset
In total, our dataset contains 625 unique queries comprising 1,326 words. As shown in Fig. 5, we report statistics including part-of-speech (POS) tag distribution, query length distribution, and the most frequent words, which collectively demonstrate the diversity of our query set across linguistic and structural dimensions.
@@ Appendix C Experiment Details and Visualization Analysis
C.1 Computational Efficiency Analysis.
Thanks to the design of the PVE module, image features are fed into the ViT only once. Subsequent iterations involve merely the SAS module and the Multi-modal Encoder, both of which are lightweight and thus enable a good trade-off between accuracy and efficiency. We analyze the computational cost and latency on a single A800 GPU, where the Baseline removes the PVE module from MSTAR.
Method GFLOPs Latency (s) FPS CTR
Baseline 248 0.0606 16.5 55.76
MSTAR 310 0.0704 14.2 60.13
Table 11: Computation and efficiency comparison on a single A800 GPU.
As shown in Table 11, MSTAR improves accuracy by 4.37% on the challenging CTR dataset, while achieving an inference speed of 14.2 FPS, over twice as fast as TG-Bridge (6.7 FPS), and maintaining comparable performance across all six benchmarks.
C.2 Text Region Localization Analysis.
We further evaluate the localization accuracy of the predicted text regions. Binary ground-truth (GT) masks are constructed by extracting the polygon coordinates of all text regions from the CTW dataset. For comparison, we also derive binary masks from the cross-attention maps of BLIP-2 using the same processing pipeline as ours. As reported in Tab. 12, we compute the Intersection over Union (IoU) between each predicted mask and its corresponding GT mask, and report both the mean IoU and the number of high-quality masks with IoU $\geq$ 0.5 across 500 images.
Method Mean IoU High-Quality Masks (IoU $\geq$ 0.5)
BLIP-2 21.78 129 (25.8%)
MSTAR 50.82 304 (60.8%)
Table 12: Quantitative comparison of text region localization on the CTW dataset. MSTAR produces substantially more accurate and higher-quality text masks than BLIP-2.
C.3 Scale-wise Evaluation on ICDAR2015.
We further evaluate MSTAR on the ICDAR2015 dataset by analyzing performance across different text scales. For each text instance, we compute the ratio between its bounding-box area and the image area using annotations from the original datasets, and group instances into scale intervals as shown in Tab  6.
For each interval, we calculate the AP score for queries whose corresponding text instances fall within the specified scale range. For instance, given the query “apple” and a target scale range $((a,b]]$, we exclude all images where “apple” appears but its area ratio is not within that range. The AP score is then computed using the remaining valid images.
Figure 6: Scale-wise performance comparison on ICDAR2015.
@@
C.4 Training Details
To ensure the reproducibility of our model, we provide the training process and the training hyper parameters, which are reported in Tab. 13. We adopt a progressive training recipe. In the first state, the model is trained on both synthetic data $D_{\text{syn}}$ and real data $D_{\text{real}}$ at an image resolution of 512. We use a learning rate of 1e-5 and linear cosine schedule with 100 warm-up steps. In this stage, we use a two-layer MLP to align the visual encoder and multi-modal encoder. In the second stage, our model is fine-tuned at the resolution of 640 only on the real data $D_{\text{real}}$. For simplicity, the hyper parameters keep the same with the first stage. In the third stage, we fine-tune the model at a higher resolution of 800 on the real data $D_{\text{real}}$. In the forth stage, the visual encoder is freezed and only the MLP, multi-modal encoder, MIM, and SAS module are optimized at a resolution of 800. For the word retrieval experiment, the synthetic data $D_{\text{syn}}$ refers to 100K images randomly sampled from SynthText-900k, and the real data $D_{\text{real}}$ comes from MLT-5K. For the multi-query experiment, $D_{\text{syn}}$ includes 50K images randomly sampled from SynthText-900k and 25k images with phrase transcriptions with the synthesis engine. $D_{\text{real}}$ is the training set from the TextCap dataset. The labels are the image captions and text transcriptions acquired with Rosetta.
C.5 Experimental Visualization Analysis
We present a qualitative analysis of the retrieval results. As illustrated in Tab.  14, our method can not only effectively leverage linguistic priors for phrase and semantic queries, but also perceive fine-grained scene text instances to achieve word and combined retrieval. For example, MSTAR successfully retrieves the combined query like “reserved” and “some” even when they appear as subtle watermarks in pictures, as depicted in the last row.
@@ C.5 Experimental Visualization Analysis
“speed limit 25” Bridge  [huang2024bridging] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
“hydrate or diedrate” is written on a water bottle Bridge  [huang2024bridging] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
“may” BLIP-2 (FT)  [li2023blip] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
“reserved” , “some” BLIP-2 (FT) [li2023blip] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Table 14: Qualitive analysis of the retrieval results. The green boxes mean correct samples and the red boxes denote false images. The queries and images are sampled from the MQTR dataset and CTW [liu2019curved] dataset.